Find the general solution and the second solution of,

x2y"-3xy'+4y=0,given y1(x)=x2



1
Expert's answer
2021-07-28T16:16:42-0400

The general solution of the differential equation will be in the form:


"y=y_1(x)+y_2(x)"

Since the solutions are linearly independent, we assume


"y_2(x) = y_1(x)v(x) \implies y_2=vx^2\qquad \cdots (1)"

Hence,


"y_2'(x) = 2vx +v'x^2 \qquad\cdots (2)\\
y_2''(x) = 2v+4xv'+x^2v'' \quad \cdots (3)"

Substitute (2) and (3) in (1):


"x^2(2v+4xv'+x^2v'')-3x(2vx+v'x^2)+4vx^2=0"


Simplifying the above:


"2vx^2+4x^3v'+x^4v''-6vx^2-3v'x^3+4vx^2=0\\
x^4v''+x^3v'=0\\"

Thus:


"xv''+v'=0 \qquad \cdots (4)"

Let:


"w=v' \qquad\cdots (5)\\
w'=v'' \qquad\cdots (6)"

Put (5) and (6) in (4)


"xw'+w=0\\
xw'=-w\\
-\frac{w'}{w}=\frac{1}{x}"

Integrate both sides


"\int-\frac{w'}{w}dw=\int \frac{1}{x}dx\\
-\ln w =\ln x\\
\ln w^{-1} = \ln x\\
w^{-1} = x \implies w=\frac{1}{x}"

Recall:


"v'(x)=w \implies v=\int w dx = \int \frac{1}{x}dx = \ln x"

Hence


"v= \ln x \qquad \cdots (7)"

Put (7) in (1), we get:


"y_2 = x^2 \ln x"

Which is the second solution.


The general solution is therefore:


"y=C_1x^2+C_2x^2 \ln x"